# A Proof of Gross’ Conjecture on $2$-Automorphic $2$-Groups

### v1

@@ 4. Finite $2$-groups with three fusion classes
In this section, we introduce the three families of AT-$2$-groups, and then prove Corollary 1.3. The first family is $A_{2}(n,\theta)$, a Suzuki $2$-group of type A, see [11, page 82].
@@
We note that, if $n=2k+1$ and $x^{\theta}=x^{k+1}$, then $A_{2}(2k+1,\theta)$ is isomorphic to a Sylow $2$-subgroup of the Suzuki simple group ${\mathrm{Sz}}(2^{2k+1})$, see [6, Proposition 13.4(vi)]. For any odd integers $n\neq 1$, the groups $A_{2}(n,\theta)$ are AT-groups by Zhang in [25, Lemma 2.2].
@@
Sylow $2$-subgroups of ${\mathrm{SU}}_{3}(2^{n})$ are Suzuki $2$-group of type B (see [12, page 314]). Equip the space ${\mathbb{F}}_{q^{2}}^{3}$ ($q=2^{n}$) with the unitary form:
@@
Then the following group $B_{2}(n)$ is a Sylow $2$-subgroup of ${\mathrm{SU}}_{3}(q)$, see [7, page 249].
@@
It was noticed in Zhang [25, Lemma 2.3] that $B_{2}(n)$ is an AT-group, with ${\mathbf{Z}}(N)=\langle M(0,b)\mid b\in{\mathbb{F}}_{q^{2}},b+b^{q}=0\rangle\cong{\mathbb{Z}}_{2}^{n}$.
@@
Let $D=\mathrm{diag}(\lambda^{-q},\lambda^{1-q},\lambda)$, where $\langle\lambda\rangle={\mathbb{F}}_{q^{2}}^{\times}$. Then $D^{-1}M(a,b)D=M(q\lambda,b\lambda^{q+1})$, and $b\lambda^{q+1}+(b\lambda^{q+1})^{q}+(a\lambda)(a\lambda)^{q}=\lambda^{1+q}(b+b^{q}+a^{1+q})=0$. Thus, $D$ induces an automorphism of $N$, denoted by $\xi$. Let $\phi$ be the Frobenius automorphism $x\mapsto x^{2}$ on ${\mathbb{F}}_{q^{2}}$, which induces an automorphism of $N$. As $M(a,b)$ and $M(q\lambda,b\lambda^{q+1})$ are not in the same coset of ${\mathbf{Z}}(N)$, it follows that $\langle\xi,\phi\rangle\cong{\Gamma\mathrm{L}}_{1}(2^{2n})$ acts faithfully on $N/{\mathbf{Z}}(N)$.
@@
For a generator $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$, define
$P(\epsilon)=\{(a,x)\in{\mathbb{F}}_{2^{6}}\times{\mathbb{F}}_{2^{3}}\mid(a,x)(b,y)=(a+b,x+y+ab^{2}\epsilon+a^{8}b^{16}\epsilon^{8})\}.$
@@
Dornhoff [8] determined finite groups $N$ of which a solvable group of automorphisms has exactly three orbits on $N$. In his classification, desired non-abelian $2$-groups are precisely $A_{2}(n,\theta)$, $B_{p}(n)$ and $P(\epsilon)$. Then the following lemma immediately holds since each Suzuki $2$-group has a solvable automorphism group by Theorem 1.1(iii).

### v2

@@ 4. Finite $2$-groups with three fusion classes
In 2020, Bors and Glasby posed a paper [2] which classified finite $2$-groups with exactly three fusion classes. In this section, we introduce the families of $2$-groups with three fusion classes, and then obtain a similar classification as in [2] by using Theorem 1.1 and the main Theorem in [9].
The first family is $A_{2}(n,\theta)$, a Suzuki $2$-group of type A, see [13, page 82].
@@
We note that, if $n=2k+1$ and $x^{\theta}=x^{k+1}$, then $A_{2}(2k+1,\theta)$ is isomorphic to a Sylow $2$-subgroup of the Suzuki simple group ${\mathrm{Sz}}(2^{2k+1})$, see [7, Proposition 13.4(vi)]. For any odd integers $n\neq 1$, the groups $A_{2}(n,\theta)$ are AT-groups by Zhang in [27, Lemma 2.2].
@@
Sylow $2$-subgroups of ${\mathrm{SU}}_{3}(2^{n})$ are Suzuki $2$-group of type B (see [14, page 314]). Equip the space ${\mathbb{F}}_{q^{2}}^{3}$ ($q=2^{n}$) with the unitary form:
@@
Then the following group $B_{2}(n)$ is a Sylow $2$-subgroup of ${\mathrm{SU}}_{3}(q)$, see [8, page 249].
@@
It was noticed in Zhang [27, Lemma 2.3] that $B_{2}(n)$ is an AT-group, with ${\mathbf{Z}}(N)=\langle M(0,b)\mid b\in{\mathbb{F}}_{q^{2}},b+b^{q}=0\rangle\cong{\mathbb{Z}}_{2}^{n}$.
@@
Let $D=\mathrm{diag}(\lambda^{-q},\lambda^{1-q},\lambda)$, where $\langle\lambda\rangle={\mathbb{F}}_{q^{2}}^{\times}$. Then $D^{-1}M(a,b)D=M(a\lambda,b\lambda^{q+1})$, and $b\lambda^{q+1}+(b\lambda^{q+1})^{q}+(a\lambda)(a\lambda)^{q}=\lambda^{1+q}(b+b^{q}+a^{1+q})=0$. Thus, $D$ induces an automorphism of $N$, denoted by $\xi$. Let $\phi$ be the Frobenius automorphism $x\mapsto x^{2}$ on ${\mathbb{F}}_{q^{2}}$, which induces an automorphism of $N$. As $M(a,b)$ and $M(a\lambda,b\lambda^{q+1})$ are not in the same coset of ${\mathbf{Z}}(N)$, it follows that $\langle\xi,\phi\rangle\cong{\Gamma\mathrm{L}}_{1}(2^{2n})$ acts faithfully on $N/{\mathbf{Z}}(N)$.
@@
For a generator $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$, define $P(\epsilon)$ to be the group consisting of elements in ${\mathbb{F}}_{2^{6}}\times{\mathbb{F}}_{2^{3}}$ with the opteration of multiplication defined by
$(a,x)(b,w)=(a+b,x+y+ab^{2}\epsilon+a^{8}b^{16}\epsilon^{8}),\mbox{ for $(a,x),(b,w)\in{\mathbb{F}}_{2^{6}}\times{\mathbb{F}}_{2^{3}}$.}$
@@
Dornhoff [9] determined finite groups $N$ of which a solvable group of automorphisms has exactly three orbits on $N$. In his classification, desired non-abelian $2$-groups are precisely $A_{2}(n,\theta)$, $B_{p}(n)$ and $P(\epsilon)$.= Then the following lemma immediately holds since each Suzuki $2$-group has a solvable automorphism group by Theorem 1.1(iii).
@@
The above lemma is the same as [2, Theorem 1.1] given by Bors and Glasby. They shows that $B_{p}(n)$ and $P(\epsilon)$ are in the family of Suzuki $2$-group of type B.